OpenStudy (anonymous):

A rectangle has width of 4 feet. Find a function that models it's area A in terms of Its perimeter P.

OpenStudy (campbell_st):

Start with what you know about the perimeter P = 2(4) + 2l or P = 2l + 8 so make l the subject. now for Area Area = l * w you know w = 4 so using what you know Area = 4l now substitute the value of l from the perimeter for the final answer

OpenStudy (anonymous):

Where do I plug that in ?

OpenStudy (campbell_st):

\[Area =4 \times l\] and the perimeter is \[P = 8 + 2l\] can you make l the subject... in the perimeter equation

OpenStudy (anonymous):

The answer is suppose to be A(P) = 2P-16 ??

OpenStudy (campbell_st):

yep... we'll get to that... in the perimeter equation above can you make length (l) the subject...